# Further results for classical and universal characters twisted by roots of unity

### v1

@@ 3. Factorization of classical characters twisted by odd powers of an even primitive root of unity
In this section, we generalize the results from our previous work [AK22] in this section. Suppose $t$ is a multiple of $4$ and $\omega$ is a primitive $t^{\prime}$th root of unity as usual. Define a tuple $X_{\omega}$ of length $tn/2+t/4$ by
$X_{\omega}=(\omega X,\omega^{3}X,\dots,\omega^{{t}-1}X,\omega,\omega^{3},\dots,\omega^{t/2-1}).$
@@ 3.2. Symplectic character factorization
We consider the symplectic character with the same specialization below.
@@
Since $\ell(\lambda)\leq\left(2n+1\right)t/4$ and $\mathrm{core}_{t/2}{(\lambda)}$ is a symplectic $(t/2)$-core, $\ell(\lambda^{(i)})+\ell(\lambda^{(t/2-2-i)})\leq 2n+1$ and by (2.17),
@@ 3.3. Even orthogonal character factorization
where the second equality comes from using 2.1. Applying it in the first column of the determinant in (3.4), we get
@@
where $W=(-X,-\overline{X},-1)$. Applying elementary column operations $C_{j}\rightarrow C_{j}+C_{j-1}$ for $j=n,\dots,2$ in order and then again using 2.1 completes the proof. ∎
We consider the even orthogonal character with the same specialization below.
@@
$\epsilon=-\sum_{i=\left\lfloor{(t+4)}/{4}\right\rfloor}^{t/2-1}\binom{n_{i}(\lambda)+1}{2}+\frac{n(n+1)}{2}+(n+1)r+\sum_{i=1}^{{(t-4)/{4}}}|{(\lambda^{(i)},-\lambda^{(t/2-i)})_{2n+1}}|.$
@@
$\epsilon=-\sum_{i=\left\lfloor\frac{t+4}{4}\right\rfloor}^{t/2-1}\binom{n_{i}(\lambda)+1}{2}+\frac{n(n+1)}{2}+(n+1)r,$
and the $(t/2)$-quotient $(\lambda^{(0)},\dots,\lambda^{(t/2-1)})$ is evaluated by considering $\lambda\in\mathcal{P}_{tn+t/2}$. Since $\ell(\lambda)\leq\left(2n+1\right)t/4$, $\mathrm{core}_{t/2}{(\lambda)}$ is an orthogonal $(t/2)$-core, $\ell(\lambda^{(i)})+\ell(\lambda^{(t/2-i)})\leq 2n+1$ and by (2.17),
@@
Now we get the required result using 3.4. ∎
@@ 3.4. Odd orthogonal character factorization
We now consider the odd orthogonal character with the same specialization.
@@
and the $(t/2)$-quotient $(\lambda^{(0)},\dots,\lambda^{(t/2-1)})$ is evaluated by considering $\lambda\in\mathcal{P}_{tn+t/2}$. Since $\ell(\lambda)\leq\left(2n+1\right)t/4$, $\mathrm{core}_{t/2}{(\lambda)}$ is a self-conjugate $(t/2)$-core, $\ell(\lambda^{(i)})+\ell(\lambda^{(t/2-1-i)})\leq 2n+1$ and by (2.7),
@@
We note that Theorems 3.3, 3.5 and 3.6 for $t=4$ can also be obtained directly from [Kum22, Theorem 2.8, Theorem 2.11 and Theorem 2.5] respectively.
@@ 4. Relations among characters of classical groups
In this section, we look at unexpected equalities between characters of different classical groups. This section is motivated by conjectures by Prasad and Wagh [PW20], which relate irreducible representations of one classical group invariant under a certain automorphism to the twisted characters of the other group. In our notation, their conjectures correspond to $t=2$. We look at more general $t$ and observe new identities for characters. We will need the following result.
@@
4.1. $\operatorname{Sp}_{2tn}$ and $\operatorname{GL}_{2tn+1}$
The first pair is the symplectic group $\operatorname{Sp}_{2tn}$ and the general linear group $\operatorname{GL}_{2tn+1}$. Recall that $\omega$ is the primitive $t$’th root of unity, $X=(x_{1},x_{2},\dots,x_{n})$ and let $\widetilde{X}=(X,\overline{X})$. Let $\mu\in\mathcal{P}_{tn}$ index a representation of $\operatorname{Sp}_{2tn}$, and construct $\lambda=(\pm\mu)_{2tn+1}$. We recall the following result.
@@ 4.1. $\operatorname{Sp}_{2tn}$ and $\operatorname{GL}_{2tn+1}$
Let $\mu\in\mathcal{P}_{tn}$ with $\mathrm{core}_{t}{(\mu)}$ symplectic, and $\lambda=(\pm\mu)_{2tn+1}$ with $\mathrm{core}_{t}{(\lambda)}=(p)$ and $p\equiv\mu_{1}\pmod{t}$. Then, for $10\leq i\leq t-$ 1, we have
@@
$\beta(\lambda)=(2\mu_{1}+2tn,\mu_{1}+\mu_{2}+2tn-1,\dots,\mu_{1}+\mu_{tn}+tn-1,\mu_{1},\\ \mu_{1}-\mu_{tn}+tn-1,\dots,\mu_{1}-\mu_{2}+1,0),$
where the first $tn$ parts are of the form $\mu_{1}+tn+1+\beta(\mu)$ and the last $tn$ parts are of the form $\mu_{1}+tn-1-\operatorname{rev}(\beta(\mu))$. Let $\beta_{j}^{(i)}(\mu)$, for $1\leq j\leq n_{i}(\mu)$, be the parts of $\beta(\mu)$ congruent to $i$ modulo $t$ arranged in decreasing order. Then
$\beta^{(i)}(\lambda)=\begin{cases}\left(\mu_{1}+tn+1+\beta^{(i-p-1)}(\mu),\,\mu_{1}+tn-1-\operatorname{rev}(\beta^{(p-i-1)}(\mu))\right)&i\neq p,\\ \left(\mu_{1}+tn+1+\beta^{(t-1)}(\mu),\,\mu_{1},\,\mu_{1}+tn-1-\operatorname{rev}(\beta^{(t-1)}(\mu))\right)&i=p.\end{cases}$
@@
Now suppose $\mathrm{core}_{t}{(\mu)}$ is symplectic. Now using Theorem 2.4 with $\widetilde{X}$ replacing $X$, $m=1$ and $y=1$, we see that $s_{\lambda}(Y,\omega Y,\dots,\allowbreak\omega^{t-1}Y,1)$ is nonzero since $\mathrm{core}_{t}{(\lambda)}$ has at most one part with value at most $t-1$. Using 4.3 and 4.1, we then obtain
@@
where $\pi_{i}=\left(\mu^{(i)},-\mu^{(t-2-i)}\right)_{2n}$ for $0\leq i\leq\left\lfloor\frac{t-3}{2}\right\rfloor$. We further factorize the Schur polynomials in the right hand side. Using Theorem 2.9 for the self-dual representation of $\operatorname{GL}_{2n}$ indexed by (±$\left(\pm\,\mu^{(t/2-1)}\right)_{2n}$, we have
@@
and Theorem 2.10 for the self-dual representation of $\operatorname{GL}_{2n+1}$ indexed by (±$\left(\pm\,\mu^{(t-1)}\right)_{2n+1}$, we have
@@
(4.2) $s_{\lambda}(\widetilde{X},\omega\widetilde{X},\dots,\omega^{t-1}\widetilde{X},1)=\operatorname{sgn}(\sigma_{\lambda}^{\mathrm{core}_{t}{(\lambda)}})\operatorname{sgn}(\sigma_{\emptyset}^{\mathrm{core}_{t}{(\lambda)}})\,\mathrm{sp}_{\mu^{(t-1)}}(X^{t})\,\mathrm{o}^{\text{even}}_{\mu^{(t-1)}}(X^{t},1)\\ \prod_{i=0}^{\left\lfloor{(t-3)}/{2}\right\rfloor}\left(s_{\pi^{(1)}_{i}}(X^{t},{\overline{X}}^{t})\right)^{2}\times\begin{cases}1&t\text{ odd,}\\ \operatorname{so}_{\mu^{(t/2-1)}}(X^{t})\,\operatorname{so}_{\mu^{(t/2-1)}}(-X^{t})&t\text{ even.}\end{cases}$
@@
This provides evidence to suggest that a further generalization of [PW20, Conjecture 3] might be true.
4.2. $\operatorname{SO}_{2tn+1}$ and $\operatorname{GL}_{2tn}$
Along similar lines as in the previous section, we let $\mu\in\mathcal{P}_{tn}$ be a partition indexing a representation of $\operatorname{SO}_{2tn+1}$, and construct $\lambda=(\pm\mu)_{2tn}$. Recall that $\mathrm{core}_{t}{(\lambda)}$ is empty if and only if $n_{i}(\lambda)=2n$ for all $i$ [AK22, Corollary 3.2].
@@ 4.2. $\operatorname{SO}_{2tn+1}$ and $\operatorname{GL}_{2tn}$
Lemma 4.6.
With $\mu$ and $\lambda$ as above, $s_{\lambda}(\widetilde{X},\omega\widetilde{X},\dots,\omega^{t-1}\widetilde{X})$ is nonzero if and only if
@@
So for all $j\in\{0,\dots,t-1\}$, the parts of $\beta(\lambda)$ less than $\mu_{1}+tn$ congruent to $j$ are in bijective correspondence with parts at least $\mu_{1}+tn$ congruent to $2p-1-j$. Hence the number of parts of $\beta(\lambda)$ at least $\mu_{1}+tn$ congruent to $j$ and $2p-1-j$ is $n_{j}(\lambda)$.
@@
With $\mu$ and $\lambda$ as above and $10\leq i\leq t-$ 1, we have
@@
Using Theorem 2.4 with $\widetilde{X}$ replacing $X$, $m=0$, we see that $s_{\lambda}(\widetilde{X},\omega\widetilde{X},\dots,\omega^{t-1}\widetilde{X})$ is nonzero since $\mathrm{core}_{t}{(\lambda)}$ is empty. Using 4.7 and 4.1,
@@
$s_{(\pm\,\mu^{({(t-1)}/{2})})_{2n}}(X^{t},\overline{X}^{t})=(-1)^{|\mu_{i}^{(t-1)/2}|}\operatorname{so}_{\mu^{({(t-1)}/{2})}}(X^{t})\operatorname{so}_{\mu^{({(t-1)}/{2})}}(-X^{t}).$
@@
$\begin{split}s_{\lambda}(\widetilde{X},\omega\widetilde{X},\dots,\omega^{t-1}\widetilde{X})=\operatorname{sgn}(\sigma_{\lambda})&\operatorname{sgn}(\sigma_{\emptyset})\,\prod_{i=0}^{\left\lfloor{(t-2)}/{2}\right\rfloor}\left(s_{(\mu^{(i)},-\mu^{(t-1-i)})_{2n}}(X^{t},{\overline{X}}^{t})\right)^{2}\\ \times&\begin{cases}\operatorname{so}_{\mu^{({(t-1)}/{2})}}(X^{t})\operatorname{so}_{\mu^{({(t-1)}/{2})}}(-X^{t})&t\text{ odd,}\\ 1&t\text{ even.}\end{cases}\end{split}$
Now, using Theorem 2.8 completes the proof. ∎
This provides evidence to suggest that a further generalization of [PW20, Conjecture 1] might be true.
4.3. $\mathrm{O}_{2tn+2}$ and $\operatorname{Sp}_{2tn}$
The third pair is the symplectic group $\operatorname{Sp}_{2tn}$ and the even orthogonal group $\mathrm{O}_{2tn+2}$. Suppose $\mu\in\mathcal{P}_{tn}$ is a partition indexing a representation of $\operatorname{Sp}_{2tn}$, and construct $\lambda=\mu_{tn}+(\mu,0)$, which has length at most $tn+1$, then $\lambda$ indexes a representation of $\mathrm{O}_{2tn+2}$.

### v2

@@ 3. Factorization of classical characters twisted by odd powers of an even primitive root of unity
We generalize the results from our previous work [AK22] in this section. Suppose $t$ is a multiple of $4$ and $\omega$ is a primitive $t^{\prime}$th root of unity as usual. Define a tuple $X_{\omega}$ of length $tn/2+t/4$ by
$X_{\omega}=(\omega X,\omega^{3}X,\dots,\omega^{{t}-1}X,\omega,\omega^{3},\dots,\omega^{t/2-1}){,}$
@@ 3.2. Symplectic character factorization
We consider the symplectic character with the same specialization below. We first recall the following result.
Lemma 3.3 ([AK22, Corollary 3.7]).
Let $\lambda$ be a partition of length at most $tn$. Then $\mathrm{core}_{t}{(\lambda)}$ is a symplectic $t$-core if and only if $n_{i}(\lambda)+n_{t-2-i}(\lambda)=2n$ for $0\leq i\leq\left\lfloor{(t-2)}/{2}\right\rfloor$ and $n_{t-1}(\lambda)=n$.
@@
Since $\ell(\lambda)\leq\left(2n+1\right)t/4$ and $\mathrm{core}_{t/2}{(\lambda)}$ is a symplectic $(t/2)$-core, $\ell(\lambda^{(i)})+\ell(\lambda^{(t/2-2-i)})\leq 2n+1$ by 3.3. Now using (2.18),
@@ 3.3. Even orthogonal character factorization
where the second equality comes from using 2.1. Applying it to the first column of the determinant in (3.4), we get
@@
where $W=(-X,-\overline{X},-1)$. On the other end, we have
$\mathrm{o}^{\text{even}}_{\lambda}(-X,-1)=\det\left(h_{\lambda_{i}-i+j}(W,-1)-h_{\lambda_{i}-i-j}(W,-1)\right).$
In this determinant, applying elementary column operations $C_{j}\rightarrow C_{j}+C_{j-1}$ for $j=n,\dots,2$ in order and then using 2.1 completes the proof. ∎
We consider the even orthogonal character with the same specialization below. First we recall the following result.
Lemma 3.6 ([AK22, Corollary 3.8]).
[…]
@@
$\epsilon=-\sum_{i=\left\lfloor{(t+4)}/{4}\right\rfloor}^{t/2-1}\binom{n_{i}(\lambda)+1}{2}+{n+1+|\lambda^{(0)}|}++\sum_{i=1}^{{(t-4)/{4}}}|{(\lambda^{(i)},-\lambda^{(t/2-i)})_{2n+1}}|.$
@@
$\epsilon=-\sum_{i=\left\lfloor{(t+4)}/{4}\right\rfloor}^{t/2-1}\binom{n_{i}(\lambda)+1}{2}+{\frac{(2n+1)(2n+2)}{2}+(2n+2)r},$
and the $(t/2)$-quotient $(\lambda^{(0)},\dots,\lambda^{(t/2-1)})$ is evaluated by considering $\lambda\in\mathcal{P}_{tn+t/2}$. Since $\ell(\lambda)\leq\left(2n+1\right)t/4$, $\mathrm{core}_{t/2}{(\lambda)}$ is an orthogonal $(t/2)$-core, $\ell(\lambda^{(i)})+\ell(\lambda^{(t/2-i)})\leq 2n+1$ by 3.6. Now using (2.18),
@@
Now we get the required result using 3.5. ∎
@@ 3.4. Odd orthogonal character factorization
We now consider the odd orthogonal character with the same specialization. Recall the following result.
Lemma 3.8 ([AK22, Corollary 3.3]).
For a partition $\lambda$ of length at most $tn$, $\mathrm{core}_{t}{(\lambda)}$ is self-conjugate if and only if
$n_{i}(\lambda)+n_{t-1-i}(\lambda)=2n,\quad 0\leq i\leq\left\lfloor{(t-1)}/{2}\right\rfloor.$
@@
and the $(t/2)$-quotient $(\lambda^{(0)},\dots,\lambda^{(t/2-1)})$ is evaluated by considering $\lambda\in\mathcal{P}_{tn+t/2}$. Since $\ell(\lambda)\leq\left(2n+1\right)t/4$, $\mathrm{core}_{t/2}{(\lambda)}$ is a self-conjugate $(t/2)$-core, $\ell(\lambda^{(i)})+\ell(\lambda^{(t/2-1-i)})\leq 2n+1$ by 3.8. Now by (2.18),
@@
We note that Theorems 3.4, 3.7 and 3.9 for $t=4$ can also be obtained directly from [Kum24, Theorem 2.8, Theorem 2.11 and Theorem 2.5] respectively.
@@ 4. Relations among characters of classical groups
In this section, we look at unexpected equalities between characters of different classical groups. This section is motivated by conjectures by Prasad and Wagh [PW22], which relate irreducible representations of one classical group invariant under a certain automorphism to the twisted characters of the other group. To be more precise, they consider the relationships between the selfdual representations of $\operatorname{SL}(2n)$ and $\text{Spin}(2n+1)$ in [PW22, Conjecture 1], and of $\operatorname{SL}(2n+1)$ and $\operatorname{Sp}(2n)$ in [PW22, Conjecture 3]. Prasad111private communication has also considered the relationships between the selfdual representations of $\operatorname{Sp}(2n)$ and $\text{Spin}(2n+2)$, as well as $E_{6}$ and $F_{4}$. We thus, compare the characters of the first three pairs of representations in this section. In our notation, their conjectures correspond to $t=1$. We look at more general $t$ and observe new identities for characters. We will need the following result.
@@
4.1. $\operatorname{Sp}({2tn})$ and $\operatorname{GL}({2tn+1})$
The first pair is the symplectic group $\operatorname{Sp}({2tn})$ and the general linear group $\operatorname{GL}({2tn+1})$. Recall that $\omega$ is the primitive $t$’th root of unity, $X=(x_{1},x_{2},\dots,x_{n})$ and let $\widetilde{X}=(X,\overline{X})$. Let $\mu\in\mathcal{P}_{tn}$ index a representation of $\operatorname{Sp}({2tn})$, and construct $\lambda=(\pm\mu)_{2tn+1}$. We recall the following result.
@@ 4.1. $\operatorname{Sp}({2tn})$ and $\operatorname{GL}({2tn+1})$
Let $\mu\in\mathcal{P}_{tn}$ with $\mathrm{core}_{t}{(\mu)}$ symplectic, and $\lambda=(\pm\mu)_{2tn+1}$ with $\mathrm{core}_{t}{(\lambda)}=(p)$ and $p\equiv\mu_{1}\pmod{t}$. Then, for $0\leq i\leq t-1$, we have
@@
$\beta(\lambda)=(2\mu_{1}+2tn,\mu_{1}+\mu_{2}+2tn-1,\dots,\mu_{1}+\mu_{tn}+tn-1,{\mu_{1}+tn},\\ \mu_{1}-\mu_{tn}+tn-1,\dots,\mu_{1}-\mu_{2}+1,0),$
where the first $tn$ parts are of the form $\mu_{1}+tn+1+\beta(\mu)$ and the last $tn$ parts are of the form $\mu_{1}+tn-1-\operatorname{rev}(\beta(\mu))$. Let $\beta^{(i)}(\mu)$ be the tuple with coordinates the parts of $\beta(\mu)$ congruent to $i$ modulo $t$ arranged in decreasing order. Then
$\beta^{(i)}(\lambda)=\begin{cases}\left(\mu_{1}+tn+1+\beta^{(i-p-1)}(\mu),\,\mu_{1}+tn-1-\operatorname{rev}(\beta^{(p-i-1)}(\mu))\right)&i\neq p,\\ \left(\mu_{1}+tn+1+\beta^{(t-1)}(\mu),\,{\mu_{1}+tn},\,\mu_{1}+tn-1-\operatorname{rev}(\beta^{(t-1)}(\mu))\right)&i=p.\end{cases}$
@@
Now suppose $\mathrm{core}_{t}{(\mu)}$ is symplectic. Now using Theorem 2.4 with $\widetilde{X}$ replacing $X$, $m=1$ and $y=1$, we see that $s_{\lambda}(Y,\omega Y,\dots,\allowbreak\omega^{t-1}Y,1)$ is nonzero since $\mathrm{core}_{t}{(\lambda)}$ has at most one part with value at most t1t-1. Using 4.3 and 4.1, we then obtain
@@
where $\pi_{i}=\left(\mu^{(i)},-\mu^{(t-2-i)}\right)_{2n}$ for $0\leq i\leq\left\lfloor\frac{t-3}{2}\right\rfloor$. We further factorize the Schur polynomials in the right hand side. Using Theorem 2.9 for the self-dual representation of $\operatorname{GL}({2n})$ indexed by $\left(\pm\,\mu^{(t/2-1)}\right)_{2n}$, we have
@@
and Theorem 2.10 for the self-dual representation of $\operatorname{GL}({2n+1})$ indexed by $\left(\pm\,\mu^{(t-1)}\right)_{2n+1}$, we have
@@
(4.2) $s_{\lambda}(\widetilde{X},\omega\widetilde{X},\dots,\omega^{t-1}\widetilde{X},1)=\operatorname{sgn}(\sigma_{\lambda}^{\mathrm{core}_{t}{(\lambda)}})\operatorname{sgn}(\sigma_{\emptyset}^{\mathrm{core}_{t}{(\lambda)}})\,\mathrm{sp}_{\mu^{(t-1)}}(X^{t})\,\mathrm{o}^{\text{even}}_{\mu^{(t-1)}}(X^{t},1)\\ \prod_{i=0}^{\left\lfloor{(t-3)}/{2}\right\rfloor}\left(s_{{\pi_{i}}}(X^{t},{\overline{X}}^{t})\right)^{2}\times\begin{cases}1&t\text{ odd,}\\ \operatorname{so}_{\mu^{(t/2-1)}}(X^{t})\,\operatorname{so}_{\mu^{(t/2-1)}}(-X^{t})&t\text{ even.}\end{cases}$
@@
This provides evidence to suggest that a further generalization of [PW22, Conjecture 3] might be true.
4.2. $\operatorname{SO}({2tn+1})$ and $\operatorname{GL}({2tn})$
Along similar lines as in the previous section, we let $\mu\in\mathcal{P}_{tn}$ be a partition indexing a representation of $\operatorname{SO}({2tn+1})$, and construct $\lambda=(\pm\mu)_{2tn}$. Recall that $\mathrm{core}_{t}{(\lambda)}$ is empty if and only if $n_{i}(\lambda)=2n$ for all $i$ [AK22, Corollary 3.12].
@@ 4.2. $\operatorname{SO}({2tn+1})$ and $\operatorname{GL}({2tn})$
So for all $j\in\{0,\dots,t-1\}$, the parts of $\beta(\lambda)$ less than $\mu_{1}+tn$ congruent to $j$ are in bijective correspondence with parts at least $\mu_{1}+tn$ congruent to 2p1j2p-1-j. Hence the number of parts of $\beta(\lambda)$ at least $\mu_{1}+tn$ congruent to $j$ and 2p1j2p-1-j is $n_{j}(\lambda)$.
@@
With $\mu$ and $\lambda$ as above and $0\leq i\leq t-1$, we have
@@
[…]
@@
$s_{(\pm\,\mu^{({(t-1)}/{2})})_{2n}}(X^{t},\overline{X}^{t})=(-1)^{|{\mu^{((t-1)/2)}}|}\operatorname{so}_{\mu^{({(t-1)}/{2})}}(X^{t})\operatorname{so}_{\mu^{({(t-1)}/{2})}}(-X^{t}).$
@@
$\begin{split}s_{\lambda}(\widetilde{X},\omega\widetilde{X},\dots,\omega^{t-1}\widetilde{X})=\operatorname{sgn}(\sigma_{\lambda})&\operatorname{sgn}(\sigma_{\emptyset})\prod_{i=0}^{\left\lfloor{(t-2)}/{2}\right\rfloor}\left(s_{(\mu^{(i)},-\mu^{(t-1-i)})_{2n}}(X^{t},{\overline{X}}^{t})\right)^{2}\\ \times&\begin{cases}{(-1)^{|{\mu^{((t-1)/2)}}|}}\operatorname{so}_{\mu^{({(t-1)}/{2})}}(X^{t})\operatorname{so}_{\mu^{({(t-1)}/{2})}}(-X^{t})&t\text{ odd,}\\ 1&t\text{ even.}\end{cases}\end{split}$
Now, using Theorem 2.8, we have that
$\begin{split}\operatorname{so}_{\mu}(X,\omega X,&\dots,\omega^{t-1}X)\;\operatorname{so}_{\mu}(-X,-\omega X,\dots,-\omega^{t-1}X)=\\ (-1)^{\epsilon}&\prod_{i=0}^{\left\lfloor{(t-2)}/{2}\right\rfloor}\left(s_{(\mu^{(i)},-\mu^{(t-1-i)})_{2n}}(X^{t},{\overline{X}}^{t})\right)^{2}\times\begin{cases}\operatorname{so}_{\mu^{({(t-1)}/{2})}}(X^{t})\operatorname{so}_{\mu^{({(t-1)}/{2})}}(-X^{t})&t\text{ odd,}\\ 1&t\text{ even.}\end{cases}\end{split}$
where
$\epsilon=\sum_{i=0}^{\left\lfloor{(t-2)}/{2}\right\rfloor}(|(\mu^{(i)}|+|\mu^{(t-1-i)}|).$
This completes the proof. ∎
This provides evidence to suggest that a further generalization of [PW22, Conjecture 1] might be true.
4.3. $\mathrm{O}({2tn+2})$ and $\operatorname{Sp}({2tn})$
The third pair is the symplectic group $\operatorname{Sp}({2tn})$ and the even orthogonal group $\mathrm{O}({2tn+2})$. Suppose $\mu\in\mathcal{P}_{tn}$ is a partition indexing a representation of $\operatorname{Sp}({2tn})$, and construct $\lambda=\mu_{tn}+(\mu,0)$, which has length at most $tn+1$, then $\lambda$ indexes a representation of $\mathrm{O}({2tn+2})$.